# Quadratic Gating Functions in Mixture of Experts: A Statistical Insight

### v2

@@ 3.1 Quadratic Polynomial Gate
$\|f_{\widehat{G}_{n}}-f_{G_{*}}\|_{L_{2}(\mu)}=\mathcal{\widetilde{O}}_{P}(n^{-1/2}).$ (13)
Proof of Theorem 3.1 is in Appendix D.1. From the result of this theorem, it can be seen that if we are able to establish the lower bound $\|f_{\widehat{G}_{n}}-f_{G_{*}}\|_{L_{2}(\mu)}\gtrsim\mathcal{L}(\widehat{G}_{n},G_{*})$ where $\mathcal{L}$ is some loss function among parameters, then we obtain the parameter estimation rate $\mathcal{L}(\widehat{G}_{n},G_{*})=\mathcal{\widetilde{O}}_{P}(n^{-1/2})$. This approach plays an vital role in establishing the rates for estimating individual parameters as well as experts in the sequel.
Turning to the parameter and expert estimation problem. A key step to establish the parameter and expert estimation rates is to decompose the discrepancy $f_{\widehat{G}_{n}}({\bm{x}})-f_{G_{*}}({\bm{x}})$ into a combination of linearly independent terms via Taylor expansions to the function $F({\bm{x}};{\bm{A}},{\bm{b}},{\bm{\eta}}):=\exp({\bm{x}}^{\top}{\bm{A}}{\bm{x}}+{\bm{b}}^{\top}{\bm{x}})h({\bm{x}},{\bm{\eta}})$. However, we notice that there is an interaction among gating parameters $A$ and $b$ expressed by the following partial differential equation (PDE):
$\dfrac{\partial F}{\partial A}({\bm{x}};{\bm{A}},{\bm{b}},{\bm{\eta}})=\dfrac{\partial^{2}F}{\partial b\partial b^{\top}}({\bm{x}};{\bm{A}},{\bm{b}},{\bm{\eta}}).$ (14)
@@
System of polynomial equations. Let $\bar{r}(m)$ be the smallest natural number $r$ such that the following system of polynomial equations does not admit any non-trivial solutions for the unknown variables: $(p_{l},\gamma_{1l},\gamma_{2l})_{l=1}^{m}\subseteq\mathbb{R}^{3}$
$\sum_{l=1}^{m}\sum_{\begin{subarray}{c}n_{1},n_{2}\in\mathbb{N}\\ n_{1}+2n_{2}=\alpha\end{subarray}}\dfrac{p_{l}^{2}~{}\gamma_{1l}^{n_{1}}~{}\gamma_{2l}^{n_{2}}}{n_{1}!~{}n_{2}!}=0,\quad\alpha=1,2,\ldots,r,$ (15)
A solution to the above system is regarded as non-trivial if all variables $p_{l}$ are non-zero, whereas at least one of the $\gamma_{1l}$ is different from zero. As shown in [Proposition 2.1, [14]], we have $\bar{r}(2)=4$, $\bar{r}(3)=6$ and $\bar{r}(m)\geq 7$ when $m\geq 4$.
Next, we introduce a condition called poly-strong identifiability to characterize the types of expert functions that admit faster estimation rates than others. From a technical view, the purpose of the strong identifiability condition is to eliminate all potential interactions among expert parameters via some PDEs as in equation (14).
@@
We say that an expert function $x\mapsto h({\bm{x}},{\bm{\eta}})$ is strongly identifiable if it is twice differentiable w.r.t its parameter ${\bm{\eta}}$, and if for any $N\geq 1$ and pair-wise different parameters ${\bm{\eta}}_{1},\ldots,{\bm{\eta}}_{N}$, the following set
@@
Example. It can be verified that the poly-strong identifiability condition holds for experts formulated as feed-forward neural networks with activation functions such as $\mathrm{ReLU}(\cdot)$ and $\tanh(\cdot)$. However, a linear expert fails to satisfy this condition.
In the sequel, we determine the parameter and expert estimation rates when using strongly identifiable experts and linear experts, respectively.
Poly-strongly identifiable experts. To capture the convergence behavior of strongly identifiable experts, let us construct a loss function among parameters based on a notion of Voronoi cells [22, 27, 26]. Given an arbitrary mixing measure $G$ with $N^{\prime}\leq N$ components, we distribute its components to the following Voronoi cells, which are generated by the components of $G_{*}$:
$\mathcal{V}_{j}\equiv\mathcal{V}_{j}(G)$ :=$:=\{i\in[N^{\prime}]:\|\omega_{i}-\omega^{*}_{j}\|\leq\|\omega_{i}-\omega^{*}_{\ell}\|,\forall\ell\neq j\},$ (16)
where $\omega_{i}:=({\bm{A}}_{i},{\bm{b}}_{i},{\bm{\eta}}_{i})$ and $\omega^{*}_{j}:=({\bm{A}}^{*}_{j},{\bm{b}}^{*}_{j},{\bm{\eta}}^{*}_{j})$ for any $j\in[N^{*}]$. Notably, the cardinality of Voronoi cell $\mathcal{V}_{j}$ is exactly the number of fitted components that approximates $\omega^{*}_{j}$. Then, the Voronoi loss function used for our analysis is given by:
$\mathcal{L}_{1}(G,G_{*}):=$ $\sum_{j:|\mathcal{V}_{j}|>1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i})\|\Delta\bm{\theta}_{ij}\|_{\frac{\bar{r}(|\mathcal{V}_{j}|)}{2},\bar{r}(|\mathcal{V}_{j}|),2}+\sum_{j:|\mathcal{V}_{j}|=1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i})\|\Delta\bm{\theta}_{ij}\|_{1,1,1}$
$+\sum_{j=1}^{N^{*}}\Big{|}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i})-\exp(c_{j}^{*})\Big{|},$ (17)
where we denote $\Delta{\bm{A}}_{ij}:={\bm{A}}_{i}-{\bm{A}}^{*}_{j}$, $\Delta{\bm{b}}_{ij}:={\bm{b}}_{i}-{\bm{b}}^{*}_{j}$, $\Delta{\bm{\eta}}_{ij}:={\bm{\eta}}_{i}-{\bm{\eta}}^{*}_{j}$, and $\Delta\bm{\theta}_{ij}=(\Delta{\bm{A}}_{ij},\Delta{\bm{b}}_{ij},\Delta{\bm{\eta}}_{ij})$. Furthermore, we define $\|\Delta\bm{\theta}_{ij}\|_{r_{1},r_{2},r_{3}}:=\|\Delta{\bm{A}}_{ij}\|^{r_{1}}+\|\Delta{\bm{b}}_{ij}\|^{r_{2}}+\|\Delta\bm{\theta}_{ij}\|^{r_{3}}$ for any $r_{1},r_{2},r_{3}\geq 1$.
Equipped with the Voronoi loss $\mathcal{L}_{1}$ defined above, we are now ready to capture the parameter estimation rate in Theorem 3.3, whose proof can be found in Appendix D.2.
@@
Suppose that the expert function $h(\cdot,{\bm{\eta}})$ is strongly identifiable, then we achieve the following lower bound for any $G\in\mathcal{G}_{N}(\Theta)$:
@@
which together with Theorem 3.1 indicates that $\mathcal{L}_{1}(\widehat{G}_{n},G_{*})=\mathcal{\widetilde{O}}_{P}(n^{-1/2})$.
@@
There are two main implications from the result of Theorem 3.3. First, it follows from the formulation of the loss function $\mathcal{L}_{1}$ that exact-specified parameters ${\bm{A}}^{*}_{j},{\bm{b}}^{*}_{j},{\bm{\eta}}^{*}_{j}$, i.e. $j\in[N^{*}]:|\mathcal{V}_{j}(\widehat{G}_{n})|=1$, share the same estimation rate of order $\widetilde{\mathcal{O}}_{P}(n^{-1/2})$. Note that as the expert $h(\cdot,{\bm{\eta}})$ is a Lipschitz function, then by denoting $\widehat{G}_{n}:=\sum_{i=1}^{\widehat{N}_{n}}\exp(\widehat{c}^{n}_{i})\delta_{(\widehat{{\bm{A}}}^{n}_{i},\widehat{{\bm{b}}}^{n}_{i},\widehat{{\bm{\eta}}}^{n}_{i})}$, we get
$\sup_{x}|h({\bm{x}},\widehat{{\bm{\eta}}}^{n}_{i})-h({\bm{x}},{\bm{\eta}}^{*}_{j})|\lesssim\|\widehat{{\bm{\eta}}}^{n}_{i}-{\bm{\eta}}^{*}_{j}\|=\widetilde{\mathcal{O}}_{P}(n^{-1/2}),$ (18)
@@
for any $i\in\mathcal{V}_{j}(\widehat{G}_{n})$. The above bound indicates that if the strongly identifiable expert $h(\cdot,{\bm{\eta}}^{*}_{j})$ is fitted by exactly one expert, it has an estimation rate is of order $\widetilde{\mathcal{O}}_{P}(n^{-1/2})$. Second, for over-specified parameters ${\bm{A}}^{*}_{j},{\bm{b}}^{*}_{j},{\bm{\eta}}^{*}_{j}$, where $j\in[N^{*}]:|\mathcal{V}_{j}(\widehat{G}_{n})|>1$, the rates for estimating them are substantially slower. In particular, the estimation rates for ${\bm{A}}^{*}_{j}$ and ${\bm{b}}^{*}_{j}$ are of orders $\widetilde{\mathcal{O}}_{P}(n^{-1/\bar{r}(|\mathcal{V}_{j}(\widehat{G}_{n})|)})$ and $\widetilde{\mathcal{O}}_{P}(n^{-1/2\bar{r}(|\mathcal{V}_{j}(\widehat{G}_{n})|)})$, respectively, which are determined by the solvability of the system (15). For instance, when those parameters are fitted by three components, the previous rates become $\widetilde{\mathcal{O}}_{P}(n^{-1/6})$ and $\widetilde{\mathcal{O}}_{P}(n^{-1/12})$. Meanwhile, parameters ${\bm{\eta}}^{*}_{j}$ enjoy an estimation rate of order $\widetilde{\mathcal{O}}_{P}(n^{-1/4})$. By arguing similarly to equation (18), the rates for estimating the experts $h(\cdot,{\bm{\eta}}^{*}_{j})$ are also $\widetilde{\mathcal{O}}_{P}(n^{-1/4})$.
Poly-weak identifiability of linear experts. We note in passing that for linear expert function $h({\bm{x}},({\bm{\beta}}_{1},\beta_{0}))=({\bm{\beta}}_{1i})^{\top}{\bm{x}}+\beta_{0i}$, where $({\bm{\beta}}_{1},\beta_{0})\in\mathbb{R}^{d}\times\mathbb{R}$, it violates the poly-strong identifiability condition due to an interaction among parameters via the following partial differential equation:
$\frac{\partial^{2}F}{\partial b\partial\beta_{0}}({\bm{x}};{\bm{A}}^{*}_{i},{\bm{b}}^{*}_{i},{\bm{\beta}}^{*}_{1i},\beta^{*}_{0i})=\frac{\partial F}{\partial{\bm{\beta}}_{1}}({\bm{x}};{\bm{A}}^{*}_{i},{\bm{b}}^{*}_{i},{\bm{\beta}}^{*}_{1i},\beta^{*}_{0i}),$ (19)
where we denote $F({\bm{x}};{\bm{A}},{\bm{b}},{\bm{\beta}}_{1},\beta_{0}):=\exp({\bm{x}}^{\top}{\bm{A}}{\bm{x}}+{\bm{b}}^{\top}{\bm{x}})({\bm{\beta}}_{1}^{\top}{\bm{x}}+\beta_{0})$. That violation leads to $\mathcal{O}(1/\log(n))$ rates of the parameters, which are considerably slower than those of strongly identifiable experts in Theorem 3.3. Please refer to Appendix C for a detailed argument of that result.
4 Practical Implications
@@
In this section, we provide two practical implications from the convergence analysis of parameter and expert estimation under the MoE models with the quadratic polynomial gate in Section 3.1 and Appendix B.
1. Benefits of quadratic monomial gating (3) over quadratic polynomial gating (2). The remarks after Theorems 3.3 and B.3 indicate that the estimation rates of the gating parameters are independent of the amount of over-specification of the number of experts and much better than those of the polynomial gating parameters, which become very slow even when we only overspecify the model by a few experts. That theoretical advantage of the monomial gating over the polynomial gating confirms the benefits of the widely used linear embeddings of the keys and queries without bias terms in the attention in practice.

### v3

@@ 3.1 Quadratic Polynomial Gate
$\|f_{\widehat{G}_{n}}-f_{G_{*}}\|_{L_{2}(\mu)}=\mathcal{O}_{P}(\sqrt{\log(n)/n}).$ (12)
The proof of Theorem 3.1 is in Appendix A.1. From the result of this theorem, it can be seen that if we are able to establish the lower bound $\|f_{\widehat{G}_{n}}-f_{G_{*}}\|_{L_{2}(\mu)}\gtrsim\mathcal{L}(\widehat{G}_{n},G_{*})$ where $\mathcal{L}$ is some loss function among parameters, then we obtain the parameter estimation rate $\mathcal{L}(\widehat{G}_{n},G_{*})=\mathcal{\widetilde{O}}_{P}(n^{-1/2})$. This approach plays an vital role in establishing the rates for estimating individual parameters as well as experts in the sequel.
Turning to the parameter and expert estimation problem. A key step to establish the parameter and expert estimation rates is to decompose the discrepancy $f_{\widehat{G}_{n}}({\bm{x}})-f_{G_{*}}({\bm{x}})$ into a combination of linearly independent terms via Taylor expansions to the function $F({\bm{x}};{\bm{A}},{\bm{b}},{\bm{\eta}}):=\exp({\bm{x}}^{\top}{\bm{A}}{\bm{x}}+{\bm{b}}^{\top}{\bm{x}})\mathcal{E}({\bm{x}},{\bm{\eta}})$. However, we notice that there is an interaction among gating parameters ${\bm{A}}$ and ${\bm{b}}$ expressed by the following partial differential equation (PDE):
$\dfrac{\partial F}{\partial{\bm{A}}}({\bm{x}};{\bm{A}},{\bm{b}},{\bm{\eta}})=\dfrac{\partial^{2}F}{\partial{\bm{b}}\partial{\bm{b}}^{\top}}({\bm{x}};{\bm{A}},{\bm{b}},{\bm{\eta}}).$ (13)
@@
System of polynomial equations. Let $\bar{r}(m)$ be the smallest natural number $r$ such that the following system of polynomial equations does not admit any non-trivial solutions for the unknown variables $(p_{l},\gamma_{1l},\gamma_{2l})_{l=1}^{m}\subseteq\mathbb{R}^{3}$:
$\sum_{l=1}^{m}\sum_{\begin{subarray}{c}n_{1},n_{2}\in\mathbb{N}\\ n_{1}+2n_{2}=\alpha\end{subarray}}\dfrac{p_{l}^{2}~{}\gamma_{1l}^{n_{1}}~{}\gamma_{2l}^{n_{2}}}{n_{1}!~{}n_{2}!}=0,\quad\alpha=1,2,\ldots,r.$ (14)
A solution to the above system is regarded as non-trivial if all variables $p_{l}$ are non-zero, whereas at least one of the $\gamma_{1l}$ is different from zero. As shown in [Proposition 2.1, Ho-Nguyen-Ann-16 ], we have $\bar{r}(2)=4$, $\bar{r}(3)=6$ and $\bar{r}(m)\geq 7$ when $m\geq 4$.
Next, we introduce a condition called poly-strong identifiability to characterize the types of expert functions that admit faster estimation rates than others. From a technical view, the purpose of the poly-strong identifiability condition is to eliminate all potential interactions among parameters expressed in the language of PDEs as in equation (13).
@@
We say that an expert function ${\bm{x}}\mapsto\mathcal{E}({\bm{x}},{\bm{\eta}})$ is strongly identifiable if it is twice differentiable w.r.t its parameter ${\bm{\eta}}$, and if for any $N\geq 1$ and distinct parameters ${\bm{\eta}}_{1},{\bm{\eta}}_{2},\ldots,{\bm{\eta}}_{N}$, the following set of functions in ${\bm{x}}$
@@
Examples. It can be verified that the poly-strong identifiability condition holds for experts formulated as two-layer FFNs with non-linear activation functions such as $\mathrm{ReLU}$ and $\mathrm{GELU}$, that is, $\mathcal{E}({\bm{x}},(\bm{\beta}_{2},\bm{\beta}_{1},\beta_{0}))=\bm{\beta}_{2}\varphi(\bm{\beta}_{1}^{\top}x+\beta_{0})$. However, linear experts of the form $\mathcal{E}({\bm{x}},(\bm{\beta}_{1},\beta_{0}))=\bm{\beta}_{1}^{\top}{\bm{x}}+\beta_{0}$ fail to satisfy this condition as the two terms $\frac{\partial\mathcal{E}}{\partial\bm{\beta}_{1}^{(u)}}({\bm{x}},(\bm{\beta}_{1},\beta_{0}))$ and ${\bm{x}}^{(u)}\cdot\frac{\partial\mathcal{E}}{\partial\beta_{0}}({\bm{x}},(\bm{\beta}_{1},\beta_{0}))$ are linearly dependent for all $u\in[d]$, that is,
$\frac{\partial\mathcal{E}}{\partial\bm{\beta}_{1}^{(u)}}({\bm{x}},(\bm{\beta}_{1},\beta_{0}))={\bm{x}}^{(u)}\cdot\frac{\partial\mathcal{E}}{\partial\beta_{0}}({\bm{x}},(\bm{\beta}_{1},\beta_{0})).$ (15)
In the sequel, we determine the parameter and expert estimation rates when using poly-strongly identifiable experts and linear experts in Section 3.1.1 and Section 3.1.2, respectively.
3.1.1 Poly-strongly Identifiable Experts
To characterize the convergence behavior of strongly identifiable experts, let us construct a loss function among parameters based on a notion of Voronoi cells manole22refined . Given an arbitrary mixing measure $G$ with $N^{\prime}\leq N$ components, we distribute its components to the following Voronoi cells, which are generated by the components of $G_{*}$:
$\mathcal{V}_{j}\equiv\mathcal{V}_{j}(G)$ :=$:=\{i\in[N^{\prime}]:\|\bm{\omega}_{i}-\bm{\omega}^{*}_{j}\|\leq\|\bm{\omega}_{i}-\bm{\omega}^{*}_{\ell}\|,\forall\ell\neq j\},$ (16)
where we denote $\bm{\omega}_{i}:=({\bm{A}}_{i},{\bm{b}}_{i},{\bm{\eta}}_{i})$ and $\bm{\omega}^{*}_{j}:=({\bm{A}}^{*}_{j},{\bm{b}}^{*}_{j},{\bm{\eta}}^{*}_{j})$ for any $j\in[N^{*}]$. Notably, the cardinality of Voronoi cell $\mathcal{V}_{j}$ is exactly the number of fitted components that approximates $\omega^{*}_{j}$. Then, the Voronoi loss function used for our analysis is given by:
$\mathcal{L}_{1}(G,G_{*}):=\sum_{j=1}^{N^{*}}\Big{|}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i})$ $-\exp(c_{j}^{*})\Big{|}+\sum_{j:|\mathcal{V}_{j}|=1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i})[\|\Delta{\bm{A}}_{ij}\|+\|\Delta{\bm{b}}_{ij}\|+\|\Delta\eta_{ij}\|]$
$+\sum_{j:|\mathcal{V}_{j}|>1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i})\Big{[}\|\Delta{\bm{A}}_{ij}\|^{\frac{\bar{r}(|\mathcal{V}_{j}|)}{2}}+\|\Delta{\bm{b}}_{ij}\|^{\bar{r}(|\mathcal{V}_{j}|)}+\|\Delta\eta_{ij}\|^{2}\Big{]},$ (17)
where we denote $\Delta{\bm{A}}_{ij}:={\bm{A}}_{i}-{\bm{A}}^{*}_{j}$, $\Delta{\bm{b}}_{ij}:={\bm{b}}_{i}-{\bm{b}}^{*}_{j}$, and $\Delta{\bm{\eta}}_{ij}:={\bm{\eta}}_{i}-{\bm{\eta}}^{*}_{j}$.
Equipped with the above Voronoi loss $\mathcal{L}_{1}$, we are now ready to capture the parameter estimation rates in Theorem 3.3, whose proof can be found in Appendix A.2.
@@ 3.1.1 Poly-strongly Identifiable Experts
Suppose that the expert function ${\bm{x}}\mapsto\mathcal{E}({\bm{x}},{\bm{\eta}})$ is poly-strongly identifiable, then we achieve the following lower bound for any $G\in\mathcal{G}_{N}(\Theta)$:
@@
which together with Theorem 3.1 indicates that $\mathcal{L}_{1}(\widehat{G}_{n},G_{*})=\mathcal{O}_{P}(\sqrt{\log(n)/n})$.
@@
There are two main implications from the results of Theorem 3.3. First, it follows from the formulation of the loss function $\mathcal{L}_{1}$ that exact-specified parameters ${\bm{A}}^{*}_{j},{\bm{b}}^{*}_{j},{\bm{\eta}}^{*}_{j}$, i.e. $j\in[N^{*}]:|\mathcal{V}_{j}(\widehat{G}_{n})|=1$, share the same estimation rate of order $\widetilde{\mathcal{O}}_{P}(n^{-1/2})$. Note that as the expert $\mathcal{E}({\bm{x}},{\bm{\eta}})$ is a Lipschitz function in ${\bm{\eta}}$, then by denoting $\widehat{G}_{n}:=\sum_{i=1}^{\widehat{N}_{n}}\exp(\widehat{c}^{n}_{i})\delta_{(\widehat{{\bm{A}}}^{n}_{i},\widehat{{\bm{b}}}^{n}_{i},\widehat{{\bm{\eta}}}^{n}_{i})}$, we get
$\sup_{{\bm{x}}}|\mathcal{E}({\bm{x}},\widehat{{\bm{\eta}}}^{n}_{i})-\mathcal{E}({\bm{x}},{\bm{\eta}}^{*}_{j})|\lesssim\|\widehat{{\bm{\eta}}}^{n}_{i}-{\bm{\eta}}^{*}_{j}\|=\widetilde{\mathcal{O}}_{P}(n^{-1/2}),$ (18)
@@
for any $i\in\mathcal{V}_{j}(\widehat{G}_{n})$. The above bound indicates that if the poly-strongly identifiable expert $\mathcal{E}({\bm{x}},{\bm{\eta}}^{*}_{j})$ is fitted by exactly one expert, it has an estimation rate of order $\widetilde{\mathcal{O}}_{P}(n^{-1/2})$. Second, for over-specified parameters ${\bm{A}}^{*}_{j},{\bm{b}}^{*}_{j},{\bm{\eta}}^{*}_{j}$, where $j\in[N^{*}]:|\mathcal{V}_{j}(\widehat{G}_{n})|>1$, the rates for estimating them are substantially slower. In particular, the estimation rates for ${\bm{A}}^{*}_{j}$ and ${\bm{b}}^{*}_{j}$ are of orders $\widetilde{\mathcal{O}}_{P}(n^{-1/\bar{r}(|\mathcal{V}_{j}(\widehat{G}_{n})|)})$ and $\widetilde{\mathcal{O}}_{P}(n^{-1/2\bar{r}(|\mathcal{V}_{j}(\widehat{G}_{n})|)})$, respectively, which are determined by the solvability of the system (14). For instance, when those parameters are fitted by three components, the previous rates become $\widetilde{\mathcal{O}}_{P}(n^{-1/6})$ and $\widetilde{\mathcal{O}}_{P}(n^{-1/12})$. Meanwhile, parameters ${\bm{\eta}}^{*}_{j}$ enjoy an estimation rate of order $\widetilde{\mathcal{O}}_{P}(n^{-1/4})$. By arguing in a similar fashion to equation (18), the rates for estimating the experts $\mathcal{E}({\bm{x}},{\bm{\eta}}^{*}_{j})$ are also $\widetilde{\mathcal{O}}_{P}(n^{-1/4})$.
3.1.2 Linear Experts
@@
Subsequently, we provide the convergence rates for parameter estimation and expert estimation under MoE with the quadratic polynomial gate.
It is worth noting that the linear expert function $\mathcal{E}({\bm{x}},({\bm{\beta}}_{1},\beta_{0}))={\bm{\beta}}_{1}^{\top}{\bm{x}}+\beta_{0i}$, where $({\bm{\beta}}_{1},\beta_{0})\in\mathbb{R}^{d}\times\mathbb{R}$, violates the poly-strong identifiability condition, leading to an interaction among parameters via the following PDE:
$\frac{\partial^{2}F}{\partial{\bm{b}}\partial\beta_{0}}({\bm{x}};{\bm{A}}^{*}_{i},{\bm{b}}^{*}_{i},{\bm{\beta}}^{*}_{1i},\beta^{*}_{0i})=\frac{\partial F}{\partial{\bm{\beta}}_{1}}({\bm{x}};{\bm{A}}^{*}_{i},{\bm{b}}^{*}_{i},{\bm{\beta}}^{*}_{1i},\beta^{*}_{0i}),$ (19)
where we denote $F({\bm{x}};{\bm{A}},{\bm{b}},{\bm{\beta}}_{1},\beta_{0}):=\exp({\bm{x}}^{\top}{\bm{A}}{\bm{x}}+{\bm{b}}^{\top}{\bm{x}})({\bm{\beta}}_{1}^{\top}{\bm{x}}+\beta_{0})$. To capture the effects of such parameter interaction on the convergence of parameter estimation, let us design another Voronoi loss tailored to this setting. More specifically, we define for any $r\geq 1$ that
Theorem 3.4.
Assume that the experts take the linear form $\mathcal{E}({\bm{x}},(\bm{\beta}_{1},\beta_{0}))={\bm{\beta}}_{1}^{\top}{\bm{x}}+\beta_{0}$, then we achieve the following minimax lower bound of estimating $G_{*}$:
$\inf_{\overline{G}_{n}\in\mathcal{M}_{N}(\Theta)}\sup_{G\in\mathcal{M}_{N}(\Theta)\setminus\mathcal{M}_{N^{*}-1}(\Theta)}\mathbb{E}_{f_{G}}[\mathcal{L}_{2,r}(\overline{G}_{n},G)]\gtrsim n^{-1/2},$
The proof of Theorem 3.4 is in Appendix A.3. A few remarks on the result of this theorem are in order. First, Theorem 3.4 reveals that using linear experts make the estimation rates for all the parameters ${\bm{A}}^{*}_{i}$, ${\bm{b}}^{*}_{i}$, ${\bm{\beta}}^{*}_{1i}$, and $\beta^{*}_{0i}$ are slower than polynomial rates $\widetilde{\mathcal{O}}_{P}(n^{-1/2r})$ for any $r\geq 1$, and, thus, could be as slow as $\mathcal{O}_{P}(1/\log^{\tau}(n))$, for some $\tau>0$, owing to the parameter interaction in equation (19). Second, we have that
Since the input space $\mathcal{X}$ is bounded, the rates for estimating linear experts $({\bm{\beta}}^{*}_{1j})^{\top}{\bm{x}}+\beta^{*}_{0j}$ could also be of order $\mathcal{O}_{P}(1/\log^{\tau}(n))$. Hence, combining with the results in Theorem 3.3, we deduce that the performance of a mixture of linear experts cannot compare to that of a mixture of non-linear experts in terms of the expert estimation problem. This observation is totally in line with the findings in chen2022theory .
3.2 Quadratic Monomial Gate
In this section, we proceed to streamline the analysis of the quadratic monomial gate based on the regression framework in equation (9). Due to the change in the gating function, the corresponding regression function is reformulated as follows:
$\tilde{f}_{G_{*}}({\bm{x}}):=\sum_{i=1}^{N^{*}}\frac{\exp({\bm{x}}^{\top}{\bm{A}}^{*}_{i}{\bm{x}}+c^{*}_{i})}{\sum_{j=1}^{N^{*}}\exp({\bm{x}}^{\top}{\bm{A}}^{*}_{j}{\bm{x}}+c^{*}_{j})}\cdot\mathcal{E}({\bm{x}},{\bm{\eta}}^{*}_{i}).$ (21)
In comparison with the quadratic polynomial gating, the first-degree monomial term ${\bm{b}}^{\top}{\bm{x}}$ has been removed from the affinity scores. As a consequence, the least squares estimator under this setting also changes accordingly to
$\widetilde{G}_{n}:=\operatorname*{arg\,min}_{G\in\mathcal{M}_{N}(\Theta)}\sum_{i=1}^{n}\Big{(}y_{i}-\tilde{f}_{G}({\bm{x}}_{i})\Big{)}^{2}.$ (22)
Given the above estimator, we provide in Theorem 3.5 the convergence rate of regression estimation $\tilde{f}_{\widetilde{G}_{n}}$ to the ground-truth regression function $\tilde{f}_{G_{*}}$.
Theorem 3.5 (Regression Estimation Rate).
$\|\tilde{f}_{\widetilde{G}_{n}}-\tilde{f}_{G_{*}}\|_{L_{2}(\mu)}=\mathcal{O}_{P}(\sqrt{\log(n)/n}).$ (23)
See Appendix A.4 for the proof of Theorem 3.5. It follows from the bound (23) that the regression estimation rate still remains parametric on the sample size, which matches that in Theorem 3.1 where we use the quadratic polynomial gate in the MoE-type regression function.
Analogous to Section 3.1, we also derive a mono-strong identifiability condition in Definition 3.6 to determine which expert functions will have faster estimation rates than others.
Definition 3.6 (Mono-strong identifiability).
We say that an expert function ${\bm{x}}\mapsto\mathcal{E}({\bm{x}},{\bm{\eta}})$ is mono-strongly identifiable if it is twice differentiable w.r.t its parameter ${\bm{\eta}}$, and if for any $k\geq 1$ and distinct ${\bm{\eta}}_{1},{\bm{\eta}}_{2},\ldots,{\bm{\eta}}_{k}$, the following set of functions in ${\bm{x}}$
[…]
is linearly independent for almost every ${\bm{x}}$.
Examples. We can validate that two-layer FFN experts with non-linear activation functions such as $\mathrm{ReLU}$ and $\mathrm{GELU}$ satisfy the mono-strong identifiability condition. On the other hand, experts of the linear form $\mathcal{E}({\bm{x}},(\bm{\beta}_{1},\beta_{0}))=\bm{\beta}_{1}^{\top}{\bm{x}}+\beta_{0}$ violate this condition.
$\mathcal{L}_{3}(G,G_{*}):=\sum_{j=1}^{N^{*}}\Big{|}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i})-\exp(c_{j}^{*})\Big{|}$ $+\sum_{j:|\mathcal{V}_{j}|=1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i})\Big{[}\|\Delta{\bm{A}}_{ij}\|+\|\Delta{\bm{\eta}}_{ij}\|\Big{]}$
$+\sum_{j:|\mathcal{V}_{j}|>1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i})\Big{[}\|\Delta{\bm{A}}_{ij}\|^{2}+\|\Delta{\bm{\eta}}_{ij}\|^{2}\Big{]}.$ (24)
Now, we are ready to capture the parameter and expert estimation rates in Theorem 3.7.
Theorem 3.7.
Assume that the expert function ${\bm{x}}\mapsto\mathcal{E}({\bm{x}},{\bm{\eta}})$ is mono-strongly identifiable, then we achieve the following lower bound for any $G\in\mathcal{M}_{N}(\Theta)$:
which together with Theorem 3.1 indicates that $\mathcal{L}_{3}(\widetilde{G}_{n},G_{*})=\mathcal{O}_{P}(\sqrt{\log(n)/n})$.
The proof of Theorem 3.7 is in Appendix A.5. A few comments regarding this theorem are in order:
(i) The rates for estimating gating parameters ${\bm{A}}^{*}_{j}$ fitted by more than one atom, i.e. $|\mathcal{V}_{j}(\widetilde{G}_{n})|>1$, are significantly improved to be of order $\widetilde{\mathcal{O}}_{P}(n^{-1/4})$. Those rates are much faster than their counterparts when using the quadratic polynomial gate, which stand at order $\widetilde{\mathcal{O}}_{P}(n^{-1/\bar{r}(|\mathcal{V}_{j}|)})$ (cf. Theorem 3.3). This rate acceleration occurs due to the disappearance of the interaction among gating parameters in equation (13). Meanwhile, the estimation rates for expert parameters ${\bm{\eta}}^{*}_{j}$ remained unchanged at the order of $\widetilde{\mathcal{O}}_{P}(n^{-1/4})$;
(ii) Model parameters ${\bm{A}}^{*}_{j},{\bm{\eta}}^{*}_{j}$ fitted by exactly one atom, i.e. $|\mathcal{V}_{j}(\widetilde{G}_{n})|=1$, enjoy the parametric estimation rates of order $\widetilde{\mathcal{O}}_{P}(n^{-1/2})$, which are comparable to their counterparts in Theorem 3.3.
Linear Experts. Similar to the case of quadratic polynomial gate, the linear expert function $\mathcal{E}({\bm{x}},({\bm{\beta}}_{1},\beta_{0}))={\bm{\beta}}_{1}^{\top}{\bm{x}}+\beta_{0}$ in the quadratic monomial gate setting does not satisfy the mono-strong idenfiability condition. That violation yields the parameter and expert estimation rates of order $\mathcal{O}(1/\log^{\tau}(n))$, for some $\tau>0$. The proof for this result is similar to that of Theorem 3.4 in Section 3.1; therefore, it is omitted here.
3.3 Practical Implications
In this section, we provide two practical implications from the convergence analysis of parameter estimation and expert estimation under the MoE models with the quadratic gating functions in Section 3.1 and Section 3.2.